# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 4 The bimodule structure and the Brauer quotients of $\mathbb{A}_{\delta}$
Assume that the block $b$ is a hyperfocal abelian Frobenius block. Then every direct summand of the $\mbox{$\mathcal{O}$}(D\times D)$-module $X$ appearing in the decomposition (4.6.1) has the form $\mathrm{Ind}_{\Delta_{\varphi}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ with $Q$ being a proper subgroup of $D$ and $\varphi:Q\longrightarrow D$ induced by an element of $E_{\mathfrak{h}}$. Moreover, there is an injective group homomorphism from $E_{\mathfrak{h}}$ to $N_{\mathbb{A}_{\delta}}(D)$ sending $\tilde{g}=gC_{G}(D)$ to $\mathfrak{e}_{\tilde{g}}$ satisfying that $(gwg^{-1})j=\mathfrak{e}_{\tilde{g}}(wj)\mathfrak{e}_{\tilde{g}}^{-1}$ for any $w\in D$. Identifying $E_{\mathfrak{h}}$ with a subgroup of $N_{\mathbb{A}_{\delta}}(D)$ through this homomorphism, then $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$ is a unitary subalgebra of $\mathbb{A}_{\delta}$ that is stable under the conjugation of $a_{u,\delta}$ and $\mathbb{A}_{\delta}=\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})\oplus Y$ as $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-$\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-bimodules with the restriction of $Y$ to $\mbox{$\mathcal{O}$}D$-$\mbox{$\mathcal{O}$}D$ being isomorphic to $X$.
@@
By Proposition 4.6, every direct summand of the $\mbox{$\mathcal{O}$}(D\times D)$-module $X$ has the form $\mbox{$\mathcal{O}$}DhD$ for some $h\in G-N_{G}(D)$. Set $Q=D\cap{{}^{h}}D$. So $Q$ is a proper subgroup of $D$. Moreover, we can assume that $Q$ is nontrivial. For in this case, we can set $\varphi=\mathrm{id}$. Let $\varphi_{h}$ be the group homomorphism from $Q$ to $D$, sending $v$ to $hvh^{-1}$. Then as $\mbox{$\mathcal{O}$}(D\times D)$-modules, $\mbox{$\mathcal{O}$}DhD$ is isomorphism to $\mathrm{Ind}_{\Delta_{\varphi_{h}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$, which is isomorphic to $\mbox{$\mathcal{O}$}D\mathop{\otimes}\limits_{\mathcal{O}Q}{{}_{\varphi_{h}}}\mbox{$\mathcal{O}$}D$. By [17, Theorem 8.7.1], we have ${{}^{h}}(Q,\bar{b}_{Q})\leq(D,\bar{b}_{D})$. Since $D$ is abelian, by [33, Theorem 2], we can assume that $h=gz$ for some $g\in N_{G}(D_{\delta})=N_{G}(D,b_{D})$ and some $z\in C_{G}(Q)$. Therefore, $\varphi_{h}=\varphi_{g}:Q\rightarrow D,~~u\mapsto gug^{-1}$. Furthermore, since $N_{G}(D_{\delta})/C_{G}(D)=PC_{G}(D)/C_{G}(D)\cdot E_{\mathfrak{h}}$, we can write $g$ as $vx$ for some $v\in P$ and some $x\in N_{G}(D_{\delta})$ with $xC_{G}(D)\in E_{\mathfrak{h}}$. By the proof of Proposition 4.6, we can deduce that $\mathbb{A}_{\delta}$ has a direct summand isomorphic to $\mathrm{Ind}_{\Delta_{\varphi_{v}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. Here, $\varphi_{v}$ is the group homomorphism from $Q$ to $D$, sending $u$ to $vuv^{-1}$. Equivalently, there is an element $\mathfrak{a}_{v}$ in $\mathbb{A}_{\delta}$ such that for any $(w,w^{\prime})\in D\times D$, $w\mathfrak{a}_{v}(w^{\prime})^{-1}=\mathfrak{a}_{v}$ if and only if $w^{\prime}=vwv^{-1}$ and $w\in Q$. At the same time, by the proof of Proposition 4.11, there is an element $y$ in $N_{G}(D_{\delta})$ such that $yC_{G}(D)$ is in $E_{\mathfrak{h}}$ and $C_{P}({{}^{y}}Q)$ is a defect group of the block $b_{{{}^{y}}Q}$ of $C_{G}({{}^{y}}Q)$. For this element $y$, there is an element $\mathfrak{a}_{y}\in\mathbb{A}_{\delta}^{*}$ such that $\mathfrak{a}_{y}(wj)\mathfrak{a}_{y}^{-1}=(ywy^{-1})j$ for any $w\in D$. Since $E_{\mathfrak{h}}$ is normal in $N_{G}(D_{\delta})/C_{G}(D)$, $(vy^{-1}v^{-1})C_{G}(D)$ still belongs to $E_{\mathfrak{h}}$. Hence, we have an element $\mathfrak{a}_{vy^{-1}v^{-1}}\in\mathbb{A}_{\delta}^{*}$ such that $\mathfrak{a}_{vy^{-1}v^{-1}}(wj)\mathfrak{a}_{vy^{-1}v^{-1}}^{-1}=(vy^{-1}v^{-1}wvyv^{-1})j$ for any $w\in D$. Then we can check that the $\mbox{$\mathcal{O}$}(D\times D)$-module generated by $\mathfrak{a}_{y}\mathfrak{a}_{v}\mathfrak{a}_{vy^{-1}v^{-1}}$ is isomorphic to $\mathrm{Ind}_{\Delta_{\varphi_{v}}({{}^{y}}Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. This implies that the $\mbox{$\mathcal{O}$}(D\times D)$-module $\mathbb{A}_{\delta}$ has a direct summand isomorphic to $\mathrm{Ind}_{\Delta_{\varphi_{v}}({{}^{y}}Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. Without loss of generality, we can assume that $C_{P}(Q)$ is a defect group of the block $b_{Q}$ of $C_{G}(Q)$.
@@
Borrowing the notation from the paragraph 3.1, there is an element $a_{v,\delta}$ in $A_{\delta}^{*}\cap\mathbb{A}v$ such that $a_{v,\delta}(uj)a_{v,\delta}^{-1}=(vuv^{-1})\cdot j$. Then $\mathfrak{a}_{v}a_{v,\delta}$ is in $j\mathbb{A}vj$ such that $w\mathfrak{a}_{v}a_{v,\delta}(w^{\prime})^{-1}=\mathfrak{a}_{v}a_{v,\delta}$ if and only if $w^{\prime}=w$ and $w\in Q$ for any $(w,w^{\prime})\in D\times D$. In particular, there is a $\mbox{$\mathcal{O}$}(D\times D)$-submodule of $j\mathbb{A}vj$ isomorphic to $\mathrm{Ind}_{\Delta(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. This implies that $(j\mathbb{A}vj)(Q)\neq 0$. By Lemma 4.10, we have $v$ being in $C_{P}(Q)$. Then the forms of indecomposable direct summand of $X$ follow from this fact.
@@
It is well-known that $N_{\mathbb{A}_{\delta}}(D)/(\mathbb{A}_{\delta}^{D})^{*}$ is isomorphic to a subgroup of $N_{A_{\delta}}(D)/(A_{\delta}^{D})^{*}$, which is isomorphic to $N_{G}(D_{\delta})/C_{G}(D)$. By the proof of Lemma 3.7, we have $N_{\mathbb{A}_{\delta}}(D)/(\mathbb{A}_{\delta}^{D})^{*}$ is in fact isomorphic to $E_{\mathfrak{h}}$. We identify these two groups with each other. At the same time, we have the following short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${k^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{\mathbb{A}_{\delta}}(D)/(j+J(\mathbb{A}_{\delta})^{D})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
@@
Since $D\rtimes E_{\mathfrak{h}}$ is a Frobenius group, we have $H^{2}(E_{\mathfrak{h}};k^{*})=1$. In particular, the short exact sequence above splits. So there is a subgroup $\hat{E}_{\mathfrak{h}}$ of $N_{\mathbb{A}_{\delta}}(D)$ containing $j+J(\mathbb{A}_{\delta})^{D}$ such that $k^{*}\cap\hat{E}_{\mathfrak{h}}=j$ and $\hat{E}_{\mathfrak{h}}/(j+J(\mathbb{A}_{\delta})^{D})\cong E_{\mathfrak{h}}$. In particular, we get another short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${j+J(\mathbb{A}_{\delta})^{D}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\hat{E}_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
@@
Since $E_{\mathfrak{h}}$ is a $p^{\prime}$-group, by [29, Lemma (45.6)], this sequence also splits. Therefore, we can view $E_{\mathfrak{h}}$ as a subgroup $\{\mathfrak{e}_{\tilde{g}}\mid\tilde{g}\in E_{\mathfrak{h}}\}$ of $N_{\mathbb{A}_{\delta}}(D)$ intersecting with $(\mathbb{A}_{\delta}^{D})^{*}$ trivially and acting on $D$ as same as $E_{\mathfrak{h}}$ acting on $D$.
@@
As $\mbox{$\mathcal{O}$}(D\times D)$-modules, by Lemma 4.5, $\mathbb{A}_{\delta}=(\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{e}_{\tilde{g}})\oplus X$ for some $\mbox{$\mathcal{O}$}(D\times D)$-submodule $X$. Clearly, $\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{e}_{\tilde{g}}$ is a unitary subalgebra of $\mathbb{A}_{\delta}$ isomorphic to $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$. Then by [17, Proposition 6.15.2], the inclusion map $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})\hookrightarrow\mathbb{A}_{\delta}$ splits as a homomorphism of $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-$\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-bimodules. Therefore, $\mathbb{A}_{\delta}=\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})\oplus Y$ for some $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-$\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-bimodule $Y$. By the Krull-Schmidt Theorem, $Y$ is isomorphic to $X$ as $\mbox{$\mathcal{O}$}D$-$\mbox{$\mathcal{O}$}D$-bimodules. ∎

### v2

@@ 4 The bimodule structure and the Brauer quotients of $\mathbb{A}_{\delta}$
Assume that the block $b$ is a hyperfocal abelian Frobenius block. Then every direct summand of the $\mbox{$\mathcal{O}$}(D\times D)$-module $X$ appearing in the decomposition (4.6.1) has the form $\mathrm{Ind}_{\Delta_{\varphi}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ with $Q$ being a proper subgroup of $D$ and $\varphi:Q\longrightarrow D$ induced by an element in $E_{\mathfrak{h}}$. Moreover, there is an injective group homomorphism from $E_{\mathfrak{h}}$ to $N_{\mathbb{A}_{\delta}}(D)$ sending $\tilde{g}=gC_{G}(D)$ to $\mathfrak{e}_{\tilde{g}}$ satisfying that $(gwg^{-1})j=\mathfrak{e}_{\tilde{g}}(wj)\mathfrak{e}_{\tilde{g}}^{-1}$ for any $w\in D$. Identifying $E_{\mathfrak{h}}$ with a subgroup of $N_{\mathbb{A}_{\delta}}(D)$ through this homomorphism, then $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$ is a unitary subalgebra of $\mathbb{A}_{\delta}$ such that $\mathbb{A}_{\delta}=\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})\oplus Y$ as $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-$\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-bimodules with $Y$ being isomorphic to $X$ as $\mbox{$\mathcal{O}$}D$-$\mbox{$\mathcal{O}$}D$-bimodules.
@@
Let $X^{\prime}$ be an indecomposable direct summand of $X$. By Proposition 4.6, we can assume that $X^{\prime}$ is isomorphic to $\mbox{$\mathcal{O}$}DhD$ for some $h\in G-N_{G}(D)$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. Set $Q={{}^{h}}D\cap D$. So $Q$ is a proper subgroup of $D$. Moreover, we can assume that $Q$ is nontrivial. For in this case, we can set $\varphi$ to be $\mathrm{id}_{D}$. As $\mbox{$\mathcal{O}$}(D\times D)$-modules, $X^{\prime}$ is isomorphism to $\mathrm{Ind}_{\Delta_{\varphi_{h}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$, which is isomorphic to $\mbox{$\mathcal{O}$}D\mathop{\otimes}\limits_{\mathcal{O}Q}{{}_{\varphi_{h}}}\mbox{$\mathcal{O}$}D$. By [16, Theorem 8.7.1], we have $(Q,\bar{b}_{Q})^{h}\leq(D,\bar{b}_{D})$. Since $D$ is abelian, by [33, Theorem 2], we can assume that $h=zg$ for some $g\in N_{G}(D_{\delta})=N_{G}(D,b_{D})$ and some $z\in C_{G}(Q)$. Therefore, $\varphi_{h}=\varphi_{g}:Q\longrightarrow D,~~u\mapsto g^{-1}ug$. Furthermore, since $N_{G}(D_{\delta})/C_{G}(D)=E_{\mathfrak{h}}\rtimes(PC_{G}(D)/C_{G}(D))$, we can write $g$ as $vx$ for some $v\in P$ and some $x\in N_{G}(D_{\delta})$ with $xC_{G}(D)\in E_{\mathfrak{h}}$. Recall from the paragraph 4.4 that there is an element $\mathfrak{a}_{x}$ in $\mathbb{A}_{\delta}^{\times}$ such that $\mathfrak{a}_{x}(uj)\mathfrak{a}_{x}^{-1}=(xux^{-1})j$ for any $u\in D$. Then $X^{\prime}\mathfrak{a}_{x}^{-1}$ is still an indecomposable direct summand of $\mathbb{A}_{\delta}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. It is routine to check that $X^{\prime}\mathfrak{a}_{x}^{-1}$ is isomorphic to ${\rm Ind}_{\Delta_{\varphi_{v}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. Equivalently, there is an element $\mathfrak{a}_{v}$ in $\mathbb{A}_{\delta}$ such that for any $(w,w^{\prime})\in D\times D$, $w\mathfrak{a}_{v}(w^{\prime})^{-1}=\mathfrak{a}_{v}$ if and only if $w^{\prime}=v^{-1}wv$ and $w\in Q$. At the same time, by the proof of Proposition 4.13, there is an element $y$ in $N_{G}(D_{\delta})$ such that $yC_{G}(D)$ is in $E_{\mathfrak{h}}$ and $C_{P}({{}^{y}}Q)$ is a defect group of the block $b_{{{}^{y}}Q}$ of $C_{G}({{}^{y}}Q)$. For this element $y$, there is an element $\mathfrak{a}_{y}\in\mathbb{A}_{\delta}^{\times}$ such that $\mathfrak{a}_{y}(wj)\mathfrak{a}_{y}^{-1}=(ywy^{-1})j$ for any $w\in D$. Since $E_{\mathfrak{h}}$ is normal in $N_{G}(D_{\delta})/C_{G}(D)$, $(v^{-1}y^{-1}v)C_{G}(D)$ still belongs to $E_{\mathfrak{h}}$. Hence, we have an element $\mathfrak{a}_{v^{-1}y^{-1}v}\in\mathbb{A}_{\delta}^{\times}$ such that $\mathfrak{a}_{v^{-1}y^{-1}v}(wj)\mathfrak{a}_{v^{-1}y^{-1}v}^{-1}=(v^{-1}y^{-1}vwv^{-1}yv)j$ for any $w\in D$. Then we can check that the $\mbox{$\mathcal{O}$}(D\times D)$-module generated by $\mathfrak{a}_{y}\mathfrak{a}_{v}\mathfrak{a}_{v^{-1}y^{-1}v}$ is isomorphic to $\mathrm{Ind}_{\Delta_{\varphi_{v}}({{}^{y}}Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. This implies that the $\mbox{$\mathcal{O}$}(D\times D)$-module $\mathbb{A}_{\delta}$ has a direct summand isomorphic to $\mathrm{Ind}_{\Delta_{\varphi_{v}}({{}^{y}}Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. Without loss of generality, we can assume that $C_{P}(Q)$ is a defect group of the block $b_{Q}$ of $C_{G}(Q)$.
@@
Borrowing the notation from the paragraph 3.1, there is an element $a_{v,\delta}$ in $A_{\delta}^{\times}\cap\mathbb{A}v$ such that $a_{v,\delta}(uj)a_{v,\delta}^{-1}=(vuv^{-1})\cdot j$. Then $\mathfrak{a}_{v}a_{v,\delta}^{-1}$ is in $j\mathbb{A}v^{-1}j$ such that $w\mathfrak{a}_{v}a_{v,\delta}^{-1}(w^{\prime})^{-1}=\mathfrak{a}_{v}a_{v,\delta}^{-1}$ if and only if $w^{\prime}=w$ and $w\in Q$ for any $(w,w^{\prime})\in D\times D$. In particular, there is an $\mbox{$\mathcal{O}$}(D\times D)$-submodule of $j\mathbb{A}v^{-1}j$ isomorphic to $\mathrm{Ind}_{\Delta(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. This implies that $(j\mathbb{A}v^{-1}j)(Q)\neq 0$. By Lemma 4.12, we have $v$ being in $C_{P}(Q)$. Therefore, the subgroup $\Delta_{\varphi_{h}}(Q)$ is equal to the subgroup $\Delta_{\varphi_{x}}(Q)$ with $xC_{G}(D)\in E_{\mathfrak{h}}$. This completes the proof of the first statement.
@@
By Proposition 3.7, $N_{\mathbb{A}_{\delta}}(D)/(\mathbb{A}_{\delta}^{D})^{\times}$ is isomorphic to $E_{\mathfrak{h}}$. We identify these two groups with each other. At the same time, we have the following short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${k^{\times}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{\mathbb{A}_{\delta}}(D)/(j+J(\mathbb{A}_{\delta}^{D}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
@@
Since $D\rtimes E_{\mathfrak{h}}$ is a Frobenius group, we have $H^{2}(E_{\mathfrak{h}};k^{\times})=1$. In particular, the short exact sequence above splits. So there is a subgroup $\hat{E}_{\mathfrak{h}}$ of $N_{\mathbb{A}_{\delta}}(D)$ containing $j+J(\mathbb{A}_{\delta}^{D})$ such that $k^{\times}\cap\hat{E}_{\mathfrak{h}}=\{j\}$ and $\hat{E}_{\mathfrak{h}}/(j+J(\mathbb{A}_{\delta})^{D})\cong E_{\mathfrak{h}}$. In particular, we get another short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${j+J(\mathbb{A}_{\delta}^{D})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\hat{E}_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
@@
Since $E_{\mathfrak{h}}$ is a $p^{\prime}$-group, by [29, Lemma (45.6)], this sequence also splits. Therefore, we can view $E_{\mathfrak{h}}$ as a subgroup $\{\mathfrak{e}_{\tilde{g}}\mid\tilde{g}\in E_{\mathfrak{h}}\}$ of $N_{\mathbb{A}_{\delta}}(D)$ intersecting with $(\mathbb{A}_{\delta}^{D})^{\times}$ trivially such that it and $E_{\mathfrak{h}}$ act in the same way on $D$.
@@
As $\mbox{$\mathcal{O}$}(D\times D)$-modules, by Lemma 4.5, $\mathbb{A}_{\delta}=(\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{e}_{\tilde{g}})\oplus X$ for some $\mbox{$\mathcal{O}$}(D\times D)$-submodule $X$. Clearly, $\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{e}_{\tilde{g}}$ is a unitary subalgebra of $\mathbb{A}_{\delta}$ isomorphic to $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$. Then by [16, Proposition 6.15.2], the inclusion map $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})\hookrightarrow\mathbb{A}_{\delta}$ splits as a homomorphism of $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-$\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-bimodules. Therefore, $\mathbb{A}_{\delta}=\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})\oplus Y$ for some $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-$\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-bimodule $Y$. By the Krull-Schmidt Theorem, $Y$ is isomorphic to $X$ as $\mbox{$\mathcal{O}$}D$-$\mbox{$\mathcal{O}$}D$-bimodules. ∎